mrs.montez is 20 years older than Mrs. cooke sixteen years ago, mrs.montez was 3 times as old as mrs. cooke was then. how old are mrs.montez and mrs. cooke now?
Answer by mananth(16946)

mrs.montez is 20 years older than Mrs. cooke. sixteen years ago mrs.montez was 3 times as old as mrs. cooke was then. how old are mrs.montez and mrs. cooke now?
let Mrs. cooke's age be x now
mrs. montez's age will be x+20
16 years ago their ages were x-16 and (x+20)-16
Mrs. montez was 3 times mrs. cooke's age then
3(x-16) = (x+20)-16
3x-48 =x+20-16
3x-x = 48+20-16
2x=52
x=26 that is cooke's age now
Mrs montez's age will be x+20 = 26+20 = 46 years